### v1

Abstract
The interplay between topology and quantum criticality gives rise to the notion of symmetry-enriched criticality, which has attracted considerable attention in recent years. However, its non-Hermitian counterpart remains largely unexplored. In this Letter, we show how parity-time ($\mathcal{PT}$) symmetry enriches non-Hermitian critical points, giving rise to a topologically distinct non-unitary universality class. By analytically investigating non-Hermitian free fermion models with $\mathcal{PT}$ symmetry, we uncover a new class of conformally invariant non-unitary critical points that host robust topological edge modes. Remarkably, the associated topological degeneracy is surprisingly encoded in the purely imaginary part of the entanglement entropy scaling—a feature absent in Hermitian systems. The underlying mechanism for the emergence of edge states at non-Hermitian criticality is traced to a generalized mass inversion that is absent in Hermitian systems.
Introduction.—Non-Hermitian (nH) quantum systems have recently attracted considerable attention due to their unique properties beyond Hermitian counterparts [1, 2, 3, 4, 5] and are closely related to diverse experimental platforms [6, 7, 8, 9, 10]. The characterization of quantum phases in nH systems is now well established [11, 12, 13, 14, 15, 16], revealing distinctive phenomena such as the non-Hermitian skin effect [17, 18, 19, 20, 21, 22] and non-Hermitian topological phases [23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35]. In contrast, phase transitions in nH systems—typically described by non-unitary conformal field theories (CFTs)—remain less understood, although they have recently drawn growing interest from the perspective of quantum entanglement [36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70].
On a different front, the universality class of quantum phase transitions can be further enriched by global symmetries, giving rise to topologically distinct universality classes, now referred to as symmetry-enriched quantum criticality (SEQC) [71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88] or, more generally, gapless symmetry-protected topological (gSPT) states  [89, 90, 91, 92, 93, 94, 95, 96, 97, 98, 99, 100, 101, 102, 103, 104, 105, 106, 107, 108, 109, 110, 111, 112, 113, 114, 115, 116, 117, 118, 119, 120, 121]. The discovery of topological physics in quantum critical systems opens new avenues for classifying phase transitions within the same universality class, fundamentally enriching the textbook understanding of phase transitions and going beyond the traditional paradigms of statistical and condensed matter physics. Remarkably, the coexistence of symmetry-protected edge states with a gapless bulk gives rise to intriguing topological phenomena absent in gapped topological phases, including nontrivial conformal boundary conditions [76, 92], algebraically localized edge modes [71, 117], universal bulk-boundary correspondence [76, 102, 83], and intrinsically gapless topological phases [96, 115].
A fundamental open question is whether the notion of SEQC can be generalized to non-unitary criticality in non-Hermitian systems, and if so, what mechanisms govern such phenomena. Despite its importance, this issue has remained largely unexplored due to the intrinsic difficulty of systematically investigating quantum criticality in generic non-Hermitian many-body systems.
To make progress in answering the above questions, we systematically explore $\mathcal{PT}$-symmetry–enriched non-unitary criticality in a family of one-dimensional nH free-fermion models, uncovering a new class of non-unitary critical points that host robust topological edge modes. Using the combination of exact solution and numerical simulations, we show that although both critical points in these models are described by non-unitary CFTs with central charge c=2c=-2, one of them is further enriched by $\mathcal{PT}$ symmetry, giving rise to topologically non-unitary criticality with stable edge states. Remarkably, the associated topological degeneracies are encoded in the purely imaginary part of the entanglement entropy scaling, a feature unique to non-Hermitian systems. Finally, we elucidate the emergence of topological edge states at this critical point by proposing an entirely new mechanism—generalized mass inversion—that is absent in Hermitian counterparts.
Non-Hermitian SSH model and its $\alpha$-extensions—We begin with the nH SSH chain with balanced gain/loss. In the momentum space, the single-particle Hamiltonian is
$\mathcal{H}_{k}=\begin{pmatrix}iu&v_{k}\\ v_{k}^{*}&-iu\end{pmatrix},\qquad v_{k}=v-we^{-ik},$ (1)
where $k$ is the single-particle momentum, $v,w\in\mathbb{R}$ are the intra-/inter-cell hoppings, and $u\in\mathbb{R}$ is the staggered imaginary on-site potential (see Fig. 1(b) for a schematic). With parity $\mathcal{P}=\sigma_{x}$ and time reversal symmetry $\mathcal{T}=\mathcal{K}$, the model satisfies $\sigma_{x}\mathcal{H}_{k}^{*}\sigma_{x}=\mathcal{H}_{k}$, i.e., it is $\mathcal{PT}$-symmetric. The bulk dispersion $E_{k}=\pm\sqrt{|v_{k}|^{2}-u^{2}}$ is real when $\min_{k}|v_{k}|>u$ ($\mathcal{PT}$-symmetric) and becomes complex otherwise, signaling spontaneous $\mathcal{PT}$ breaking.
From the bulk-boundary correspondence, the number of edge-mode pairs under the open boundary condition (OBC) is directly related to the winding number computed in the momentum space [28]
$\omega=\frac{1}{2\pi}\int_{-\pi}^{\pi}\!dk\,\arg v_{k},$ (2)
which depends only on $v_{k}$. The vertical dashed line $w=v$ in Fig. 1(a) therefore marks a change in edge-mode count that is independent of $u$. Note that inside the $\mathcal{PT}$-broken wedge the dashed line is not a phase boundary, because the Berry (Zak) phase, which determines the phase structure, no longer corresponds to the winding number there [122].
Turning on $u$ has a genuinely nH effect on criticality. At $u=0$ the Hermitian SSH chain has a single quantum critical point (QCP) at $w=v$ (black dot). For $u>0$, this point splits into two straight critical lines $u=\pm(w-v)$ that border the $\mathcal{PT}$-broken wedge. The left critical line lies in the $\omega=0$ sector and is topologically trivial. The right critical line lies in $\omega=1$; it is topological and is the main focus of this work. We then extend the construction to models with richer topological content in this work.
To realize a broader class of nH QCPs with nontrivial topology, we generalize the nH SSH model using an $\alpha$-chain construction inspired by Hermitian counterparts [72]. Specifically, we define the $\alpha$-nH SSH model as a $\mathcal{PT}$-symmetric free-fermion chain with long-range hopping,
$v_{k}^{(\alpha)}=v\,e^{-i(\alpha-1)k}-w\,e^{-i\alpha k}.$ (3)
The case $\alpha=1$ recovers the nH SSH model. For $\alpha\geq 2$, long-range hopping preserves the three-region structure of Fig. 1(a) while increasing the edge-mode multiplicity. In the $\mathcal{PT}$-symmetric gapped regime |vw|>u|v-w|>u under OBC, the side $v>w$ hosts $\alpha-1$ edge-mode pairs (winding number $\omega=\alpha-1$), whereas the side $w>v$ hosts $\alpha$ pairs (winding number $\omega=\alpha$). Accordingly, there are two nH critical lines that lie in distinct winding sectors: the one with lower winding number ($\omega=\alpha-1$) and the one with higher winding number ($\omega=\alpha$). In the main text, we focus on the $\alpha=1$ case and include the $\alpha=2$ model as a representative higher $\alpha$ example (see Fig. 1(c) for a schematic).
Refer to caption(a)
Refer to caption(b)
Refer to caption(c)
Figure 1: (a) Phase diagram of the $\mathcal{PT}$-symmetric nH SSH models in the (u,wv)(u,w-v) plane. The Hermitian critical point at $(0,0)$ splits into two non-Hermitian critical lines $u=\pm(w-v)$ that border the $\mathcal{PT}$-broken region. The dashed line $w=v$ separates winding sectors $\omega=0$ and $\omega=1$. (b) Schematic nH SSH chain: red/blue circles denote $\pm iu$; black/purple bonds denote Hermitian hoppings. (c) Schematic $\alpha=2$ generalization.
Detecting non-unitary CFT at non-Hermitian criticality.—We assess conformal criticality in the $\alpha$-nH SSH models via entanglement scaling and a complementary Casimir check. Using the correlation matrix method [122], the entropy for a subsystem of length $\ell_{A}$ is
$S_{A}=\sum_{n}\left[-\nu_{n}\log\nu_{n}-(1-\nu_{n})\log(1-\nu_{n})\right],$ (4)
[…]
First we review the ordinary ($\alpha=1$) nH SSH model. At its trivial QCP, the correlation eigenvalues are real and occur in $(\nu,1{-}\nu)$ pairs, ensuring that the entanglement entropy remains real and scales with central charge c=2c=-2, consistent with a non-unitary CFT (see Fig. 2(a)).
In contrast, the topologically nontrivial QCP exhibits a complex conjugate pair of correlation eigenvalues, $\nu_{\pm}=1/2\pm i\mathcal{I}$, which lead to an imaginary contribution to the entanglement entropy. While previous studies [36] encountered ambiguities due to the multivalued complex logarithm, we restore the correct analytic structure by consistently shifting one logarithmic branch by $2\pi$. This yields an additional term $i\pi$ in the entropy, while preserving the scaling form with c=2c=-2 in the real part (see Fig. 2(b)). Unlike in Hermitian systems, the entropy here acquires a constant imaginary part; we defer its physical interpretation to a later section. Full details of the branch-cut prescription are provided in the Supplementary Material [122].
We now turn to the $\alpha=2$ model to examine how these features generalize. At the QCP with higher winding number ($\omega=2$), the entanglement spectrum contains two complex-conjugate eigenvalue pairs. Applying the same branch-cut prescription, the entropy acquires an additional term $2i\pi$, while maintaining the c=2c=-2 scaling (see Fig. 2(c)). The doubled imaginary part reflects the contribution from both complex pairs.
A complementary signature of criticality comes from the finite-size scaling of the ground-state energy under periodic boundary condition (PBC). In CFT, the ground-state energy takes the form
$E_{0}^{\mathrm{PBC}}(L)=L\epsilon-\frac{\pi v_{F}c}{6L}+\cdots,$ (5)
where $\epsilon$ is the bulk energy density and $v_{F}$ is the Fermi velocity.
For all $\alpha$-nH SSH models, the bulk dispersion remains identical across phases, enabling a unified scaling analysis near the critical points. As shown in Fig. 2(d), the finite-size corrections to the ground-state energy exhibit clear $1/L$ behavior with a coefficient consistent with c=2c=-2, corroborating the entanglement entropy results. Additional results for OBC and higher-$\alpha$ models—covering both entanglement entropy and ground-state energy—are fully consistent with c=2c=-2 and presented in the SM. Together, these findings demonstrate that all critical points in the $\alpha$-nH SSH models are described by a non-unitary CFT with central charge c=2c=-2.
Refer to caption(a)
Refer to caption(b)
Refer to caption(c)
Refer to caption(d)
Figure 2: Entanglement entropy $S_{A}$ as a function of the subsystem size $\ell_{A}$ for a total system length $L=10000$. Blue dots indicate the real part of $S_{A}$, red dots indicate the imaginary part, and the blue line shows the numerical fit. (a) $\alpha=1$, $(v,w,u)=(2,1,1-10^{-12})$ — trivial QCP, $\operatorname{Re}[S_{A}]=-0.665\ln[\sin(\pi\ell_{A}/L)]-10.18$. (b) $\alpha=1$, $(v,w,u)=(1,2,1-10^{-12})$ — topological QCP, $\operatorname{Re}[S_{A}]=-0.665\ln[\sin(\pi\ell_{A}/L)]-10.87$. (c) $\alpha=2$, $(v,w,u)=(1,2,1-10^{-12})$ — higher winding number QCP, $\operatorname{Re}[S_{A}]=-0.665\ln[\sin(\pi\ell_{A}/L)]-11.35$. (d) Finite-size scaling of the ground-state energy in the $\alpha$-nH SSH model under PBC with $v_{F}=\sqrt{2}$. All results support that these QCPs are described by a non-unitary CFT with central charge c=2c=-2.
Entanglement–edge correspondence and the imaginary entropy—At the critical points of the $\alpha$-nH SSH chain, we observe a correspondence between the physical and entanglement spectra. Under OBC, the system supports localized edge modes with purely imaginary energies $E=\pm iu$. Under PBC, the entanglement spectrum features complex-conjugate eigenvalue pairs of the form $\nu=1/2\pm i\mathcal{I}$, whose number matches that of the OBC edge modes [see Fig. 3(a,b)]. Using the standard correlation–entanglement mapping $\nu=(1+e^{\epsilon})^{-1}$, equivalently
$\epsilon=\log\frac{1-\nu}{\nu}=\pm 2i\arctan(2\mathcal{I}),$ (6)
corresponding to a purely imaginary entanglement energy. This generalizes the Li–Haldane picture: although both energy are no longer real or degenerate, the one-to-one counting and the vanishing real parts ($\operatorname{Re}[E]=\operatorname{Re}[\epsilon]=0$) still match.
Refer to caption(a)
Refer to caption(b)
Figure 3: Entanglement spectrum of the $\alpha$–nH SSH model in the topological phase. The left panels show the real part $\operatorname{Re}[\nu]$ and the right panels show the imaginary part $\operatorname{Im}[\nu]$ of the eigenvalues. (a) For $\alpha=1$, a single complex-conjugate pair is present. (b) For $\alpha=2$, two complex-conjugate pairs appear. In both cases, the number of complex-conjugate pairs matches the number of edge modes.
A key consequence concerns the imaginary part of the entanglement entropy. Requiring the correct c=2c=-2 scaling at all critical points forces us to adopt the same branch-cut convention used in the previous section. With this convention, each complex-conjugate pair contributes $i\pi$ to $S_{A}$. Together with the entanglement spectrum–edge correspondence, at a QCP with winding number $\omega$,
$S_{A}=-\frac{2}{3}\log\ell_{A}-i\pi\omega+\mathrm{const.}$ (7)
This identifies a subleading imaginary contribution $i\pi\omega$. Unlike in Hermitian SEQC, where topological features at criticality are reflected only in the entanglement spectrum [101, 83], the nH case renders them directly visible in the imaginary part of the entanglement entropy—a feature unique to nH systems.
It is then natural to extend the same branch choice into the adjacent SPT phase by continuity. With that extension, $\operatorname{Im}[S_{A}]$ takes the same constant value throughout each fixed-$\omega$ sector. Thus $\operatorname{Im}S_{A}$ serves as a topological index reflecting the winding number $\omega$ across the $\mathcal{PT}$-symmetric sector.
$\mathcal{PT}$ symmetry-protected topological edge states at non-unitary criticality.—The $\alpha$-nH SSH chain supports gapless topological features at criticality. We now show that these features are protected by $\mathcal{PT}$ symmetry, and that the protection is lost once $\mathcal{PT}$ symmetry is broken. From the dispersion $E_{k}=\pm\sqrt{|v_{k}|^{2}-u^{2}}$ the $\mathcal{PT}$-symmetric regime is characterized by the requirement $|v_{k}|\geq u$ for all $k$. In particular, this enforces $v_{k}\neq 0$ throughout the Brillouin zone, so the loop traced by $v_{k}$ never crosses the origin and the winding number $\omega$ is pinned. Equivalently, the number of OBC edge-mode pairs is fixed within a $\mathcal{PT}$-symmetric sector; this description includes the QCPs themselves.
[…]
To further test the stability of the edge structure, we add moderate disorder to both the hopping amplitudes and the on-site potentials while preserving global $\mathcal{PT}$ symmetry and keeping the system at criticality. As shown in Fig. 4(a), the imaginary part of the entanglement entropy, $\operatorname{Im}[S_{A}]$, remains the same across all disorder realizations. This indicates that the topological structure is robust against disorder.
Together, these results demonstrate that the generalized $\alpha$-nH SSH model hosts a robust topological edge states protected by $\mathcal{PT}$ symmetry, independent of a bulk gap or translational invariance. The critical point thus realizes a $\mathcal{PT}$ symmetry-enriched non-unitary CFT.
Refer to caption(a)
Refer to caption(b)
Figure 4: (a) Disorder-averaged entanglement entropy $S_{A}$ at its topological QCP ($v(x)=1+\delta(x)$, $w=2$, $u(x)=1-\delta(x)-10^{-10}$; $L=1000$; $\delta(x)\in[-0.999,0.999]$). Averaged over 10001000 disorder realizations; error bars denote ±1 s.e.m. across realizations. Blue dots represent the real part of $S_{A}$, red dots represent the imaginary part. (b) Interface schematic between a trivial gapped region (left) and the nH SSH chain at its topological QCP (right). Shown are the mass profile $m(x)$ (black), the spectral gap $\Delta(x)=\sqrt{m(x)^{2}-u(x)^{2}}$ (blue), and the bound-state density $|\psi(x)|^{2}$ (red) localized at the interface.
Generalized mass inversion in $\mathcal{PT}$ symmetry-enriched non-Hermitian QCP—A distinctive feature of the $\alpha$-nH SSH model is the appearance of an intermediate $\mathcal{PT}$-broken symmetry-breaking phase absent in Hermitian free-fermion systems. This additional phase shifts the location of the critical point and is closely tied to the persistence of edge modes across the transition. In the Hermitian $\alpha\geq 2$-chain, topological edge states rely on kinetic inversion with long-range hopping to maintain edge states at criticality. By contrast, in the nH SSH model, the survival of edge modes follows from a simple generalization of mass inversion without requiring long-range hopping ($\alpha=1$).
Near the critical point, the low-energy continuum Hamiltonian takes the Dirac-like form
$\mathcal{H}=\begin{bmatrix}iu&-\partial_{x}+m\\ \partial_{x}+m&-iu\end{bmatrix}$, (8)
where $m$ is the mass profile and $u$ the non-Hermitian strength. The eigenvalue equations
$(-\partial_{x}+m)\psi_{B}(x)$ $=(E-iu)\psi_{A}(x),$ (9)
$(\partial_{x}+m)\psi_{A}(x)$ $=(E+iu)\psi_{B}(x)$ (10)
decouple at $E=\pm iu$, yielding localized edge solutions with decay length $\sim 1/m$ for $m>0$ even as the bulk gap $\sqrt{m^{2}-u^{2}}$ closes.
More generally, this is due to when we have more than one mass operators $\{m_{a}\Gamma_{a}\}$ with $\Gamma_{a}$ being the gamma matrices in the Dirac-like Hamiltonian. Here some of the $m_{a}$ can be imaginary. Edge localization is set by the domain-wall mass—the component that flips across the boundary—so $\xi\sim 1/|m_{bdy}|$. The bulk gap, however, is controlled by the spectral mass $\Delta=\sqrt{\sum_{a}m_{a}^{2}}$. Thus these scales are able to decouple when we have multiple mass term. Including this scale decoupling together with ordinary mass inversion is what we refer to as generalized mass inversion.
This mechanism is absent in Hermitian systems, where the gap is always enhanced and cannot close without delocalizing the edge modes. This makes a free-fermion SEQC without long-range hopping possible only in the nH setting. For higher $\alpha$, a generalized mass inversion coexists with kinetic inversion, see the SM [122].
The schematic in Fig. 4(b) depicts an interface geometry (left: trivial gapped phase; right: nH SSH at its topological QCP). Unlike the OBC case, the interface equations remain coupled and lack a simple closed-form solution; nevertheless, generalized mass inversion pins an edge state at the interface [122].
Discussion and concluding remarks.—To summarize, we uncover a new class of non-unitary criticality enriched by $\mathcal{PT}$ symmetry, which hosts robust topological edge states. Specifically, we construct a broad family of one-dimensional $\mathcal{PT}$-symmetric nH free-fermion models with $\alpha$-range hopping, all exhibiting critical points described by non-unitary CFTs with central charge c=2c=-2, as confirmed numerically through the scaling of entanglement entropy and ground-state energy. More importantly, we unambiguously demonstrate that these nH critical points support nontrivial edge states protected by $\mathcal{PT}$ symmetry. The associated topological degeneracy is reflected in the bulk entanglement spectrum—representing a generalized Li-Haldane correspondence at nH criticality—and, remarkably, is also encoded in the imaginary part of the entanglement entropy, a feature unique to nH systems. Finally, we analytically show that the topological edge modes at these QCPs are stabilized by a new mechanism, generalized mass inversion—without requiring long-range hopping—highlighting a distinct protection mechanism from Hermitian counterparts.
Looking ahead, identifying symmetry-enriched nH criticalities in higher dimensions and in interacting many-body systems represents an interesting direction for future work. Regarding experimental realization, we note that nH free-fermion lattice models and entanglement-based observables can be implemented and probed using quantum and classical simulators, including ultracold atoms [123, 124], superconducting quantum processors [125, 126], and phononic systems [127]. Our work not only generalizes the notion of topology to the realm of non-unitary criticality but also contributes to a deeper understanding of critical phenomena in nH systems.
Note added: After completing this manuscript, we became aware of a related independent study on the critical edge states in one-dimensional nH free-fermion chains [128].
Acknowledgement: We thank Xueda Wen for helpful discussion. K.-H. Chou and P.-Y. Chang was supported by National Science and Technology Council of Taiwan under Grants No. NSTC 113-2112-M- 007-019, 114-2918-I-007-015. X.-J. Yu was supported by the National Natural Science Foundation of China (Grant No.12405034) and a start-up grant from Fuzhou University. P.-Y.C acknowledges support from RIKEN Center for Interdisciplinary Theoretical and Mathematical Sciences and National Center for Theoretical Sciences, Physics Division. This research was supported in part by Perimeter Institute for Theoretical Physics. Research at Perimeter Institute is supported by the Government of Canada through the Department of Innovation, Science, and Economic Development, and by the Province of Ontario through the Ministry of Colleges and Universities.